# Shelf-Supervised Multi-Modal Pre-Training for 3D Object Detection

### v1

@@ Abstract
State-of-the-art 3D object detectors are often trained on massive labeled datasets. However, annotating 3D bounding boxes remains prohibitively expensive and time-consuming, particularly for LiDAR. Instead, recent works demonstrate that self-supervised pre-training with unlabeled data can improve detection accuracy with limited labels. Contemporary methods adapt best-practices for self-supervised learning from the image domain to point clouds (such as contrastive learning). However, publicly available 3D datasets are considerably smaller and less diverse than those used for image-based self-supervised learning, limiting their effectiveness. We do note, however, that such data is naturally collected in a multimodal fashion, often paired with images. Rather than pre-training with only self-supervised objectives, we argue that it is better to bootstrap point cloud representations using image-based foundation models trained on internet-scale image data. Specifically, we propose a shelf-supervised approach (e.g. supervised with off-the-shelf image foundation models) for generating zero-shot 3D bounding boxes from paired RGB and LiDAR data. Pre-training 3D detectors with such pseudo-labels yields significantly better semi-supervised detection accuracy than prior self-supervised pretext tasks. Importantly, we show that image-based shelf-supervision is helpful for training LiDAR-only and multi-modal (RGB + LiDAR) detectors. We demonstrate the effectiveness of our approach on nuScenes and WOD, significantly improving over prior work in limited data settings.
Keywords: Shelf-Supervised 3D Object Detection, Vision-Language Models, Autonomous Vehicles
@@ 1 Introduction
Bootstrapping 3D Representations with 2D Priors. VLMs [15, 16, 17, 18] are trained on internet-scale data and show impressive zero-shot detection performance across many application domains [19]. Therefore, we posit that pre-training with noisy pseudo-labels distilled from vision-language models can outperform contrastive features learned from scratch (cf. Fig. 1). However, distilling image-based foundational knowledge to LiDAR remains an open challenge [20, 21, 22]. We address this challenge by explicitly “inflating” 2D instance masks to 3D bounding boxes via unlabelled multimodal (RGB + LiDAR) training data, using a careful combination of LiDAR cues, HD maps, and shape priors [10]. This allows us to produce accurate 3D pseudo-labels which can then be used to pre-train uni-modal (LiDAR) or multi-modal (RGB + LiDAR) or 3D detectors. One significant advantage of our shelf-supervision over self-supervision is that it naturally produces a more aligned pretext task; instead of pre-training 3D detectors with contrastive learning, we learn from 3D pseudo-labels.
@@ 2 Related Works
Self-Supervised Learning for 3D Representations has been applied to many domains, including object-centric point clouds [41, 42, 43], indoor scenes [44, 45, 46, 47, 48, 49], and outdoor environments [9, 7, 50]. Current techniques for self-supervised point cloud pre-training can be broadly characterized by their receptive field. Scene-based contrastive learning [6, 51] (often used in the context of representation learning for indoor scenes) may not capture the necessary fine-grained details required for recognizing small objects. In contrast, voxel-based contrastive learning (often used in the context of representation learning for outdoor scenes) inherently struggles with a restricted receptive field, limiting its ability to encode larger structures. Region-based representations [7, 52] attempt to strike a balance between coarse-grained scene representations and fine-grained voxel representations. However, most regions in outdoor scenes (e.g. ground plane and buildings) lack informative cues, making it difficult to learn robust features. More recently, multi-modal contrastive learning methods [8, 9] demonstrate that paired multi-modal sensors can provide complementary (self-)supervision for pre-training.
PointContrast [6] established correspondences between different views of point clouds using a contrastive loss. DepthContrast [14] treated different depth maps as instances and discriminates between them. STRL [53] extracted invariant representations from temporally correlated frames in a 3D point cloud sequence. More recently, SimIPU [8] and CALICO [9] delve into multi-modal contrastive learning by utilizing paired RGB and LiDAR data. GeoMAE [50] frames point-cloud representation learning as a masked auto-encoding task, and uses geometric pretext tasks including occupancy, normals, and curvature estimation. In contrast to prior work, we argue that directly pre-training with noisy pseudo-labels from vision foundation models can be a more effective pre-training strategy for 3D object detection.
@@ 3.1 Generating 3D Bounding Box Pseudo-Labels
2D Mask Generation. The first stage of our pipeline requires producing accurate 2D instance masks for a fixed vocabulary of object categories. We prompt a VLM detector (e.g. Detic [15] or GroundingDINO [18]) with class names (e.g. car, bus, truck) to generate 2D box proposals. Though some VLM detectors (i.e., Detic) already produce instance segmentation masks, we found it more effective to prompt separate (foundational) models such as SAM [57] with the predicted 2D bounding boxes to generate high-quality segmentation masks. We evaluate the impact of different VLMs on pseudo-label quality in the supplement.
@@
3D Cuboid Generation. We define an initial candidate 3D centroid to be the medoid of points within each mask, expressed as $m=\operatorname*{\arg\!\min}_{x\in P_{M}}\sum_{p\in P_{M}}{||p-x||_{2}}$, where $P_{M}$ is the set of LiDAR points that lie within the mask $M$. Note that this medoid tends to lie on the surface of the object visible to the LiDAR sensor rather than the true object centroid, which implies it will require refinement. To estimate a canonical width-height-length, we simply prompt an LLM (ChatGPT) with each class name and use the returned value for all instances. Although LLMs do not have access to our 3D training data, they have seen many descriptions of object shapes on the web and can provide reasonable prototypical 3D sizes of common objects. Lastly, we estimate box orientation for vehicles using lane geometry from an HD map and assign the direction of the nearest lane to each cuboid. We assign an orientation of zero degrees to all non-vehicles (e.g. pedestrians, barriers and traffic cones).
3D Refinement. Fig. 2 visualizes our overall pseudo-label generation pipeline. Because many components of our pipeline provide only rough 3D estimates (e.g., not all trucks have precisely the same 3D dimension), we find that 3D labels can be refined via self-training, which is discussed in the next section. The supplement ablates several strategies for improvement, including prompt engineering to improve VLM zero-shot 2D detections, LiDAR accumulation to improve medoid estimation, mask erosion to remove noisy LiDAR points near mask boundaries, medoid compensation to reduce object center estimation biases, and late-fusion of independent zero-shot 3D detectors to improve orientation and shape estimation. The supplement also discusses limitations of our pseudo-label method. However, despite these limitations, we find that imperfect pseudo-labels still provide an effective signal for pre-training.
@@ 3.2 Training with Pseudo-Labels
In this paper, we evaluate CM3D by training two popular detectors, CenterPoint [62] and BEVFusion [63]. Following CALICO’s [9] experiment setup, we train BEVFusion with the CenterHead and turn off class-balanced grouping and sampling. Although copy-paste augmentation is widely used when training 3D detectors, we find that performance degrades with our noisy pseudo-labels. We train the CenterHead following standard 3D detection losses. We use the sigmoid focal loss (for recognition) [64] and L1 regression loss (for localization). Concretely, our loss function is as follows: $L=L_{HM}+\lambda L_{REG}$, where $L_{HM}=\sum_{i=0}^{C}SigmoidFocalLoss(X_{i},Y_{i})$ and $L_{REG}=|X_{BOX}-Y_{BOX}|$, where $X_{i}$ and $Y_{i}$ are the $i^{th}$ class’ predicted and ground-truth heat maps, while $X_{BOX}$ and $Y_{BOX}$ are the predicted and pseudo-label box attributes.
Self Training. Although directly pre-training with our pseudo-labels significantly improves mAP, it also yields lower NDS compared to prior work. We posit that this is due to the errors in size and orientation estimation, which can both be attributed to our simplistic 3D priors. Prior work suggests that self-training can be effective for semi-supervised learning [35, 21]; not only can it discover more objects [35, 21], but it may also refine noisy box labels. We find that self-training significantly improves orientation and shape estimates by exploiting the regularity of multimodal data; similar-looking objects will generate similar pseudo-labels due to the smoothness of the learned detection function. We ablate the impact of self-training in the supplement.
@@ 4.1 Experiment Setup
Datasets. We evaluate our work using the well-established nuScenes dataset [1]. We follow the suggested protocol in [7, 50, 9] and sample Kof the training data uniformly from the entire training set. Lastly, we also evaluate our approach on the Waymo Open Dataset  [2] in the supplement, and find that our conclusions hold across datasets.
@@ 4.2 Comparison to State-of-the-Art
We first evaluate the zero-shot performance of CM3D to highlight the quality of our pseudo-labels. Next, we evaluate the performance of detectors trained on pseudo-labels. Lastly, we evaluate the performance of detectors pre-trained on CM3D psuedo-labels and fine-tuned on labeled data. Note that models are pre-trained with pseudo-labels from (1 - K%) of the training set and fine-tuned with ground truth labels from Kof the training set. We copy the results of prior work from CALICO and refer readers to [9] for further details. We present salient conclusions from Table 1 below.
@@
Distilling Multi-Modal Priors into Uni-Modal Models. Although CM3D requires paired RGB images and LiDAR sweeps for pseudo-label generation, we can easily train RGB-only or LiDAR-only student models (e.g. CenterPoint + CM3D). Importantly, these student models don’t require paired RGB-LiDAR data at inference (and are therefore compared against other models that only use LiDAR at inference). However, we find that the CenterPoint + CM3D (0% Training Data) student model performs worse than the pseudo-label generator (CM3D). This can be attributed to learning with noisy labels and not having access to the same sensor data as the teacher. However, we find that using multi-modal cues for both training and testing yields higher results. Specifically, BEVFusion + CM3D (0% Training Data) achieves 3.9 mAP and 1.9 NDS better performance than CenterPoint + CM3D (0% Training Data).
@@
      CenterPoint [62] + Rand. Init.   L   L   33.1   37.4
@@
      BEVFusion [63] + Rand. Init.   L + C   L + C   39.0   43.7
@@
      CenterPoint [62] + Rand. Init.   L   L   41.1   48.0
@@
      BEVFusion [63] + Rand. Init.   L + C   L + C   46.2   51.9
@@
      CenterPoint [62]+ Rand. Init.   L   L   47.1   56.7
@@
      BEVFusion [63] + Rand. Init.   L + C   L + C   53.1   58.9
@@
      CenterPoint [62] + Rand. Init.   L   L   53.2   61.0
@@
      BEVFusion [63] + Rand. Init.   L + C   L + C   58.5   61.8
@@
Significant Improvements in Low-Data Setting. Pre-training CenterPoint with CM3D pseudo-labels and fine-tuning on labeled data improves over the prior art of PRC [9] by 8.1 mAP and 2.8 NDS with 5% labeled data. Similarly, pre-training BEVFusion with CM3D pseudo-labels outperforms prior art (CALICO) by 8.6 mAP and 4.6 NDS. This suggests that aligning your pre-training and fine-tuning tasks can yield better performance.
@@
Qualitative Results. We visualize the output of our pseudo-label generation approach in Fig. 4. Although our method generates reasonable predictions in many cases, we find that our method fails in cases of occlusions (where there is no 3D information) and in cases where the vision-language model predicts a false positive detection with high confidence. See Fig 4 for more detailed analysis.
@@ 5 Conclusion
In this paper, we propose Cross-Modal 3D Detection Distillation (CM3D), a zero-shot approach for generating 3D bounding boxes using vision-language models. We demonstrate that pre-training detectors with zero-shot 3D bounding box pseudo-labels achieves higher downstream semi-supervised detection than prior contrastive learning objectives. However, we find that our proposed pre-training strategy limits generalization to arbitrary downstream tasks. We conduct comprehensive experiments to ablate our design choices and demonstrate that our simple method significantly improves over prior work in limited data settings.
@@ Appendix A CM3D Pseudo-Label Refinement
When training all detectors with pseudo-labels, we employ standard augmentation techniques (except for copy-paste augmentation, see Appendix C). Following established practices, we aggregate the past 10 sweeps for LiDAR densification using the provided ego-vehicle poses. We train CenterPoint and BEVFusion using the same hyperparameters prescribed by their respective papers. For CenterPoint, we first train the detector for 20 epochs with CM3D pseudo-labels and fine-tune the detector for 20 epochs using the limited set of annotations. For BEVFusion, we first pre-train the LiDAR-only branch using CM3D pseudo-labels for 20 epochs and fine-tune the LiDAR-only branch for 20 epochs using the limited set of annotations. We train the fusion model (RGB + LiDAR) for 6 epochs using the limited amount of labeled training data. Lastly, we fine-tune all models using self-training. Specifically, we use the fine-tuned model to generate new pseudo-labels on the unlabeled portion of the train-set and retrain the detector on the entire train-set (including the limited set of ground truth labels and pseudo-labels) from scratch. We ablate self-training further in Appendix B. We conduct all experiments on 8 RTX 3090 GPUs.

### v2

# Shelf-Supervised Cross-Modal Pre-Training for 3D Object Detection

@@ Abstract
State-of-the-art 3D object detectors are often trained on massive labeled datasets. However, annotating 3D bounding boxes remains prohibitively expensive and time-consuming, particularly for LiDAR. Instead, recent works demonstrate that self-supervised pre-training with unlabeled data can improve detection accuracy with limited labels. Contemporary methods adapt best-practices for self-supervised learning from the image domain to point clouds (such as contrastive learning). However, publicly available 3D datasets are considerably smaller and less diverse than those used for image-based self-supervised learning, limiting their effectiveness. We do note, however, that such data is naturally collected in a multimodal fashion, often paired with images. Rather than pre-training with only self-supervised objectives, we argue that it is better to bootstrap point cloud representations using image-based foundation models trained on internet-scale image data. Specifically, we propose a shelf-supervised approach (e.g. supervised with off-the-shelf image foundation models) for generating zero-shot 3D bounding boxes from paired RGB and LiDAR data. Pre-training 3D detectors with such pseudo-labels yields significantly better semi-supervised detection accuracy than prior self-supervised pretext tasks. Importantly, we show that image-based shelf-supervision is helpful for training LiDAR-only and multi-modal (RGB + LiDAR) detectors. We demonstrate the effectiveness of our approach on nuScenes and WOD, significantly improving over prior work in limited data settings. Our code is available on GitHub.
Keywords: Shelf-Supervised 3D Object Detection, Semi-Supervised Learning, Vision-Language Models, Autonomous Vehicles
@@ 1 Introduction
Bootstrapping 3D Representations with 2D Priors. VLMs [15, 16, 17] are trained on internet-scale data and show impressive zero-shot detection performance across many application domains [19]. Therefore, we posit that pre-training with noisy pseudo-labels distilled from vision-language models can outperform contrastive features learned from scratch (cf. Fig. 1). However, distilling image-based foundational knowledge to LiDAR remains an open challenge [20, 21, 22]. We address this challenge by explicitly “inflating” 2D instance masks to 3D bounding boxes via unlabelled multimodal (RGB + LiDAR) training data, using a careful combination of LiDAR cues, HD maps, and shape priors [10]. This allows us to produce accurate 3D pseudo-labels which can then be used to pre-train uni-modal or multi-modal 3D detectors. One significant advantage of our shelf-supervision over self-supervision is that it naturally produces a more aligned pretext task; instead of pre-training 3D detectors with contrastive learning, we learn from 3D pseudo-labels.
@@ 2 Related Works
Self-Supervised Learning for 3D Representations has been applied to many domains, including object-centric point clouds [41, 42, 43], indoor scenes [44, 45, 46, 47, 48, 49], and outdoor environments [9, 7, 50]. Current techniques for self-supervised point cloud pre-training can be broadly characterized by their receptive field. Scene-based contrastive learning [6, 51] (often used in the context of representation learning for indoor scenes) may not capture the necessary fine-grained details required for recognizing small objects. In contrast, voxel-based contrastive learning (often used in the context of representation learning for outdoor scenes) inherently struggles with a restricted receptive field, limiting its ability to encode larger structures. Region-based representations [7, 52] attempt to strike a balance between coarse-grained scene representations and fine-grained voxel representations. However, most regions in outdoor scenes (e.g. ground plane and buildings) lack informative cues, making it difficult to learn robust features. More recently, multi-modal contrastive learning methods [8, 9] demonstrate that paired multi-modal sensors can provide complementary (self-)supervision for pre-training. PointContrast [6] established correspondences between different views of point clouds using a contrastive loss. DepthContrast [14] treated different depth maps as instances and discriminates between them. STRL [53] extracted invariant representations from temporally correlated frames in a 3D point cloud sequence. More recently, SimIPU [8] and CALICO [9] delve into multi-modal contrastive learning by utilizing paired RGB and LiDAR data. GeoMAE [50] frames point-cloud representation learning as a masked auto-encoding task, and uses geometric pretext tasks including occupancy, normals, and curvature estimation. In contrast to prior work, we argue that directly pre-training with noisy pseudo-labels from vision foundation models can be a more effective pre-training strategy for 3D object detection.
@@ 3.1 Generating 3D Bounding Box Pseudo-Labels
2D Mask Generation. The first stage of our pipeline requires producing accurate 2D instance masks for a fixed vocabulary of object categories. We prompt a VLM detector (e.g. Detic [15] or GroundingDINO [18]) with class names (e.g. car, bus, truck) to generate 2D box proposals. Though some VLM detectors (i.e., Detic) already produce instance segmentation masks, we found it more effective to prompt separate (foundational) models such as SAM [57] with the predicted 2D bounding boxes to generate high-quality segmentation masks. We evaluate the impact of using different VLMs in Appendix F.
@@
3D Cuboid Generation. We define an initial candidate 3D centroid to be the medoid of points within each mask, expressed as $m=\operatorname*{\arg\!\min}_{x\in P_{M}}\sum_{p\in P_{M}}{||p-x||_{2}}$, where $P_{M}$ is the set of LiDAR points that lie within the mask $M$. Note that this medoid tends to lie on the surface of the object visible to the LiDAR sensor rather than the true object centroid, which implies it will require refinement. To estimate a canonical width-height-length, we simply prompt an LLM (ChatGPT) with each class name and use the returned value for all instances. Although LLMs do not have access to our 3D training data, they have seen many descriptions of object shapes on the web and can provide reasonable prototypical 3D sizes of common objects (Appendix G). Lastly, we estimate box orientation for vehicles using lane geometry from an HD map and assign the direction of the nearest lane to each cuboid. We assign an orientation of zero degrees to all non-vehicles (e.g. pedestrians, barriers and traffic cones).
3D Refinement. Fig. 2 visualizes our overall pseudo-label generation pipeline. Because many components of our pipeline provide only rough 3D estimates (e.g., not all trucks have precisely the same 3D dimension), we find that 3D labels can be refined via self-training, which is discussed in the next section. Appendix A ablates several strategies for improvement, including prompt engineering to improve VLM zero-shot 2D detections, LiDAR accumulation to improve medoid estimation, mask erosion to remove noisy LiDAR points near mask boundaries, medoid compensation to reduce object center estimation biases, and late-fusion of independent zero-shot 3D detectors to improve orientation and shape estimation. We find that imperfect pseudo-labels still provide an effective signal for pre-training.
@@ 3.2 Training with Pseudo-Labels
In this paper, we evaluate CM3D by training two popular detectors, CenterPoint [62] and BEVFusion [63]. Following CALICO’s [9] experiment setup, we train BEVFusion with the CenterHead and turn off class-balanced grouping and sampling. Although copy-paste augmentation is widely used when training 3D detectors, we find that performance degrades with our noisy pseudo-labels (Appendix D). We train the CenterHead following standard 3D detection losses. We use the sigmoid focal loss (for recognition) [64] and L1 regression loss (for localization). Concretely, our loss function is as follows: $L=L_{HM}+\lambda L_{REG}$, where $L_{HM}=\sum_{i=0}^{C}SigmoidFocalLoss(X_{i},Y_{i})$ and $L_{REG}=|X_{BOX}-Y_{BOX}|$, where $X_{i}$ and $Y_{i}$ are the $i^{th}$ class’ predicted and ground-truth heat maps, while $X_{BOX}$ and $Y_{BOX}$ are the predicted and pseudo-label box attributes.
Self Training. Although directly pre-training with our pseudo-labels significantly improves mAP, it also yields lower NDS compared to prior work. We posit that this is due to the errors in size and orientation estimation, which can both be attributed to our simplistic 3D priors. Prior work suggests that self-training can be effective for semi-supervised learning [35, 21]; not only can it discover more objects [35, 21], but it may also refine noisy box labels. We find that self-training significantly improves orientation and shape estimates by exploiting the regularity of multimodal data; similar-looking objects will generate similar pseudo-labels due to the smoothness of the learned detection function. We ablate the impact of self-training in the Appendix B.
@@ 4.1 Experiment Setup
Datasets. We evaluate our work using the well-established nuScenes dataset [1]. We follow the suggested protocol in [7, 50, 9] and sample Kof the training data uniformly from the entire training set. Lastly, we also evaluate our approach on the Waymo Open Dataset  [2] in Appendix I, and find that our conclusions hold across datasets.
@@ 4.2 Comparison to State-of-the-Art
We first evaluate the class-based zero-shot performance of CM3D to highlight the quality of our pseudo-labels. We benchmark class-agnostic performance in Appendix E. Next, we evaluate the performance of detectors trained on pseudo-labels. Lastly, we evaluate the performance of detectors pre-trained on CM3D psuedo-labels and fine-tuned on labeled data. Note that models are pre-trained with pseudo-labels from the entire training set and fine-tuned with ground truth labels from Kof the training set. We copy the results of prior work from CALICO and refer readers to [9] for further details. We present salient conclusions from Table 1 below.
@@
Distilling Multi-Modal Priors into Uni-Modal Models. Although CM3D requires paired RGB images and LiDAR sweeps for pseudo-label generation, we can easily train RGB-only or LiDAR-only student models (e.g. CenterPoint + CM3D). Importantly, these student models don’t require paired RGB-LiDAR data at inference (and are therefore compared against other models that only use LiDAR at inference). However, we find that the CenterPoint + CM3D (0% Training Data) student model performs worse than the pseudo-label generator (CM3D). This can be attributed to learning with noisy labels and not having access to the same sensor data as the teacher. Interestingly, we find that using multi-modal cues for both training and testing yields higher results.
@@
      CenterPoint [62] (Rand. Init.)   L   L   33.1   37.4
@@
      BEVFusion [63] (Rand. Init.)   L + C   L + C   39.0   43.7
@@
      CenterPoint [62] (Rand. Init.)   L   L   41.1   48.0
@@
      BEVFusion [63] (Rand. Init.)   L + C   L + C   46.2   51.9
@@
      CenterPoint [62] (Rand. Init.)   L   L   47.1   56.7
@@
      BEVFusion [63] (Rand. Init.)   L + C   L + C   53.1   58.9
@@
      CenterPoint [62] (Rand. Init.)   L   L   53.2   61.0
@@
      BEVFusion [63] (Rand. Init.)   L + C   L + C   58.5   61.8
@@
Significant Improvements in Low-Data Setting. Pre-training CenterPoint with CM3D pseudo-labels and fine-tuning on labeled data improves over the prior art of PRC [9] by 8.1 mAP and 2.8 NDS with 5% labeled data. Similarly, pre-training BEVFusion with CM3D pseudo-labels outperforms prior art (CALICO) by 8.6 mAP and 4.6 NDS. Notably all pre-training strategies improve over random initialization (Rand. Init.). This suggests that aligning your pre-training and fine-tuning tasks can yield better performance.
@@
Qualitative Results. We visualize the output of our pseudo-label generation approach in Fig. 4. Although our method generates reasonable predictions in many cases, we find that our method fails in cases of occlusions (where there is no 3D information) and in cases where the VLM predicts a false positive detection with high confidence. See Fig 4 and Appendix H for detailed analysis.
@@ 5 Conclusion
In this paper, we propose Cross-Modal 3D Detection Distillation (CM3D), a zero-shot approach for generating 3D bounding boxes using vision-language models. We demonstrate that pre-training detectors with zero-shot 3D bounding box pseudo-labels achieves higher downstream semi-supervised detection than prior contrastive learning objectives. We conduct comprehensive experiments to ablate our design choices and demonstrate that our simple method significantly improves over prior work in limited data settings. However, we find that our proposed pre-training strategy limits generalization to arbitrary downstream tasks. Moreover, our current pseudo-labeling pipeline requires LiDAR for precise box localization and HD maps for orientation estimation, which may not be readily available for broader robotics domains. We expand on the limitations and opportunities for future work in Appendix J.
6 Acknowledgements
We would like to thank Jenny Seidenschwarz for her feedback on early drafts. This work was supported in part by NSF GRFP (Grant No. DGE2140739).
@@ Appendix A CM3D Pseudo-Label Refinement
When training all detectors with pseudo-labels, we employ standard augmentation techniques (except for copy-paste augmentation, see Appendix D). Following established practices, we aggregate the past 10 sweeps for LiDAR densification using the provided ego-vehicle poses. We train CenterPoint and BEVFusion using the same hyperparameters prescribed by their respective papers. For CenterPoint, we first train the detector for 20 epochs with CM3D pseudo-labels and fine-tune the detector for 20 epochs using the limited set of annotations. For BEVFusion, we first pre-train the LiDAR-only branch using CM3D pseudo-labels for 20 epochs and fine-tune the LiDAR-only branch for 20 epochs using the limited set of annotations. We train the fusion model (RGB + LiDAR) for 6 epochs using the limited amount of labeled training data. Lastly, we fine-tune all models using self-training. Specifically, we use the fine-tuned model to generate new pseudo-labels on the unlabeled portion of the train-set and retrain the detector on the entire train-set (including the limited set of ground truth labels and pseudo-labels) from scratch. We ablate self-training further in Appendix B. We conduct all experiments on 8 RTX 3090 GPUs.